# Diffusion Models as Network Optimizers: Explorations and Analysis

### v3

@@ III-C Training and Sampling
Input: Dataset $\mathcal{D}$ with paired samples $(x,y)$, Number of diffusion steps $T$, Noise schedule factors $(\alpha_{t})_{t=1}^{T}$, Probability of unconditional sample $p_{uncond}$, Model parameters $\theta$;
@@
3       Sample a data instance $(x,y)\in\mathcal{D}$;
@@ 4       Sample time step $t\sim\text{Uniform}(1,T)$;
5       Sample noise $\epsilon\sim\mathcal{N}(0,I)$;
6       Compute noisy data $y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$;
7       Set the condition input $x=\varnothing$ with probability $p_{uncond}$;
8       Compute model’s predicted noise $\epsilon_{\theta}(y_{t},x,t)$;
@@ 9       Compute the loss:
$\mathcal{L}_{\theta}=\|\epsilon-\epsilon_{\theta}(y_{t},x,t)\|^{2}$
@@
Input: Number of diffusion steps $T$, Trained model $\theta$, Noise schedule factors $(\alpha_{t})_{t=1}^{T}$, Input parameter $x$;
Output: Generated solution $y_{0}$;
@@
2Sample initial noise $y_{T}\sim\mathcal{N}(0,I)$;
@@
4       Sample $\epsilon\sim\mathcal{N}(0,I)$ if $t>1$, else set $z=0$;
5       Update $y_{t-1}$ as:
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\epsilon}_{\theta}(y_{t},x,t)\right)+\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\epsilon$
@@ III-C1 Learning Objective and Loss Definition
In the ideal scenario, we hope to use maximum likelihood estimation to make $D_{KL}(p(y_{0})||p_{\theta}(y_{0}))\approx 0$, that is, the difference between the target distribution $p(y_{0})$ and the distribution $p_{\theta}(y_{0})$ generated by the model is close to zero. However, since it is computationally challenging to directly solve $p(y_{0})$ and perform full trajectory integration, it must consider all possible reverse trajectories to obtain a more accurate estimate. In practical applications, alternative objective functions such as denoising score matching, $\epsilon$-prediction, and the Evidence Lower Bound (ELBO) are commonly employed [66]. We adopt the $\epsilon$-prediction strategy to approximate $\nabla_{y}\log p(y_{0}|y_{t},x,t)$, thereby capturing gradient information across varying noise levels. The model learns to recover high-quality results from noisy data by minimizing the following loss function
$\mathcal{L}_{\theta}=\mathbb{E}_{y,x,\epsilon\sim\mathcal{N}(0,\boldsymbol{I}),t}[\lVert\epsilon-\epsilon_{\theta}(y_{t},t,x)\rVert^{2}].$ (14)
Eq. 14 is iteratively used during training by repeatedly uniformly sampling $t$. The detailed process of training is outlined in Algorithm 1.
@@ IV-A2 Datasets
We construct our dataset based on the problem formulations from previous studies [9, 44, 5]. The dataset consists of paired samples in the form $(x,y)$, where $x$ represents the input variables and $y$ denotes the corresponding optimal solution. To obtain the optimal solutions, we apply an exhaustive search method. Specifically, for the CO problem, we conduct experiments involving joint optimization of offloading decisions and resource allocation in a setup with one edge server and three terminal nodes. For the MSR problem, we evaluate cases with $3$ and $80$ channels, with total available power of $10$W and $20$W, respectively. In the NU problem, we use a scenario involving three terminal nodes and one UAV node. The training dataset sizes for these three problems are $30,000$ samples for CO, $10,000$ samples for each of the MSR cases, and $10,000$ samples for NU.
@@ IV-A5 Metrics
We set the ratio of the objective function value corresponding to the output solution $y_{0}$ and the ground truth $y_{true}$ as the evaluation criterion, i.e. $\frac{f(x,y_{0})}{f(x,y_{true})}$, named Exceed_ratio. The closer the ratio is to 1, the better.
@@ IV-B2 Strength of Condition
IV-B3 Number of Diffusion Steps
For the number of diffusion steps $T$. As shown in Fig. 5, $T$ that is too large or too small will both make poor learning results. We believe that it is because $T$ is too small to ensure the Gaussian result distribution of the forward process causing wrong reverse process, while $T$ that is too large will bring more knowledge to be learned, resulting in insufficient training epochs and model capacity. Large number of diffusion steps may also result in redundant denoising costs.
@@ IV-C Denoising Process
Figure 8: Objective values of the denoising process for 5 different samples of the CO problem and their corresponding optimal solution benchmarks. For example, $x_{1}$ represents a feature input, with its solution generation process depicted using a solid line. $y_{1}^{*}$ denotes the ground truth solution, represented by a dashed line indicating its corresponding optimal cost.
Figure 9: Objective values of the denoising process for 5 different samples of the MSR problem and their corresponding optimal solution benchmarks. The meaning of $x$ and $y$ are consistent with those in Fig. 8.
Figure 10: Objective values of the denoising process for 5 different samples of the NU problem and their corresponding optimal solution benchmarks. The meaning of $x$ and $y$ are consistent with those in Fig. 8.
@@ IV-C1 CO Problem
As shown in Fig. 8, the horizontal axis represents the denoising time steps, while the vertical axis shows the global cost of computation offloading, which we aim to minimize. Each colored line corresponds to the denoising process of a sample’s solution, with each point representing the global cost at a specific time step. The dashed lines indicate the optimal total cost for each input’s optimal solution; for example, the red dashed line marks the optimal cost of the blue line’s input $x_{1}$ and its optimal solution $y_{1}^{*}$.
@@ IV-C2 MSR Problem
Similarly, in Fig. 9, the vertical axis shows the total transmission rate, which we aim to maximize. Other settings of Fig. 9 follow the same format as Fig. 8. Since the high-quality solution space of individual inputs to this problem occupies a large area of the feasible solution space, the initial noise solution has performance close to the optimal solution, such as $x_{2}$ in Fig. 9. In addition, we observe that MSR converges faster than CO, which aligns with the fact that MSR, as a convex optimization problem, has a lower complexity than the MINLP required for CO.
@@ V Conclusion
In this work, we bridge the theoretical gaps in GDM as an emerging network optimizer and provide further empirical guidance. This novel network optimization approach holds the potential to deliver significant improvements in both efficiency and quality for IoT networks. Our study has provided a concise theoretical demonstration of the inherent advantages of generative models over discriminative models in optimization solving. Specifically, GDMs offer a promising approach by learning the underlying high-quality solution distribution of feasible space and enabling repeated sampling from this distribution, which improves their ability to handle complex solution spaces and achieve optimal results. By capturing global solution structures rather than directly mapping inputs to outputs, GDMs address the inherent limitations of discriminative models. Our work has leveraged these strengths of GDMs, demonstrating their capacity to not only enhance the flexibility and accuracy of network optimization but also streamline the optimization process across various high-dimensional and complex scenarios.

### v5

@@ III-C Training and Sampling
Input: Dataset $\mathcal{D}$ with paired samples $(\mathbf{x},\mathbf{y})$, Number of diffusion steps $T$, Noise schedule factors $(\alpha_{t})_{t=1}^{T}$, Probability of unconditional sample $p_{uncond}$, Model parameters $\theta$;
@@
3       Sample a data instance $(\mathbf{x},\mathbf{y})\in\mathcal{D}$;
@@ 4       Sample time step $t\sim\text{Uniform}(1,T)$;
5       Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$;
6       Compute noisy data $\mathbf{y}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$;
7       Set the condition input $\mathbf{x}=\varnothing$ with probability $p_{uncond}$;
8       Compute model’s predicted noise $\boldsymbol{\epsilon}_{\theta}(\mathbf{y}_{t},\mathbf{x},t)$;
@@ 9       Compute the loss:
$\mathcal{L}_{\theta}=\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{y}_{t},\mathbf{x},t)\|^{2}$
@@
Input: Number of diffusion steps $T$, Trained model $\theta$, Noise schedule factors $(\alpha_{t})_{t=1}^{T}$, Input parameter $\mathbf{x}$;
Output: Generated solution $\mathbf{y}_{0}$;
@@
2Sample initial noise $\mathbf{y}_{T}\sim\mathcal{N}(0,\mathbf{I})$;
@@
4       Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else set $\boldsymbol{\epsilon}=0$;
5       Update $\mathbf{y}_{t-1}$ as:
$\mathbf{y}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{y}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{y}_{t},\mathbf{x},t)\right)+\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
@@ III-C1 Learning Objective and Loss Definition
In the ideal scenario, we hope to use maximum likelihood estimation to make $D_{KL}(p(\mathbf{y}_{0})||p_{\theta}(\mathbf{y}_{0}))\approx 0$, that is, the difference between the target distribution $p(\mathbf{y}_{0})$ and the distribution $p_{\theta}(\mathbf{y}_{0})$ generated by the model is close to zero. However, since it is computationally challenging to directly solve $p(\mathbf{y}_{0})$ and perform full trajectory integration, it must consider all possible reverse trajectories to obtain a more accurate estimate. In practical applications, alternative objective functions such as denoising score matching, $\boldsymbol{\epsilon}$-prediction, and the Evidence Lower Bound (ELBO) are commonly employed [65]. We adopt the $\boldsymbol{\epsilon}$-prediction strategy to approximate $\nabla_{\mathbf{y}}\log p(\mathbf{y}_{0}|\mathbf{y}_{t},\mathbf{x},t)$, thereby capturing gradient information across varying noise levels. The model learns to recover high-quality results from noisy data by minimizing the following loss function
$\mathcal{L}_{\theta}=\mathbb{E}_{\mathbf{y},\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),t}[\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{y}_{t},t,\mathbf{x})\rVert^{2}].$ (14)
Eq. 14 is iteratively used during training by repeatedly uniformly sampling $t$, detailed in Algorithm 1.
@@ IV-A2 Datasets
We construct our dataset based on the problem formulations from previous studies [8, 43, 4]. The dataset consists of paired samples in the form $(\mathbf{x},\mathbf{y})$, where $\mathbf{x}$ represents the input variables and $\mathbf{y}$ denotes the corresponding optimal solution. To obtain the optimal solutions, we apply an exhaustive search method. Specifically, for the CO problem, we conduct experiments involving joint optimization of offloading decisions and resource allocation in a setup with one edge server and three terminal nodes. For the MSR problem, we evaluate cases with $3$ and $80$ channels, with total available power of $10$W and $20$W, respectively. In the NU problem, we use a scenario involving three terminal nodes and one UAV node. The training dataset sizes for these three problems are $30,000$ samples for CO, $10,000$ samples for each of the MSR cases, and $10,000$ samples for NU.
@@ IV-A5 Metrics
We set the ratio of the objective function value corresponding to the output solution $\mathbf{y}_{0}$ and the ground truth $\mathbf{y}_{true}$ as the evaluation criterion, i.e. $\frac{f(\mathbf{x},\mathbf{y}_{0})}{f(\mathbf{x},\mathbf{y}_{true})}$, named Exceed_ratio. The closer the ratio is to 1, the better.
@@ IV-B2 Strength of Condition
IV-B3 Number of Diffusion Steps
For the number of diffusion steps $T$. As shown in Fig. 5, $T$ that is too large or too small will both make poor learning results. We believe that it is because $T$ is too small to ensure the Gaussian result distribution of the forward process causing the wrong reverse process, while $T$ that is too large will bring more knowledge to be learned, resulting in insufficient training epochs and model capacity. A large number of diffusion steps may also result in redundant denoising costs.
@@ IV-C Denoising Process
Figure 8: Objective values of the denoising process for 5 different samples of the CO problem and their corresponding optimal solution benchmarks. For example, $\mathbf{x}_{1}$ represents a feature input, with its solution generation process depicted using a solid line. $\mathbf{y}_{1}^{*}$ denotes the ground truth solution, represented by a dashed line indicating its corresponding optimal cost.
Figure 9: Objective values of the denoising process for 5 different samples of the MSR problem and their corresponding optimal solution benchmarks. The meaning of $\mathbf{x}$ and $\mathbf{y}$ are consistent with those in Fig. 8.
Figure 10: Objective values of the denoising process for 5 different samples of the NU problem and their corresponding optimal solution benchmarks. The meaning of $\mathbf{x}$ and $\mathbf{y}$ are consistent with those in Fig. 8.
@@ IV-C1 CO Problem
As shown in Fig. 8, the horizontal axis represents the denoising time steps, while the vertical axis shows the global cost of computation offloading, which we aim to minimize. Each colored line corresponds to the denoising process of a sample’s solution, with each point representing the global cost at a specific time step. The dashed lines indicate the optimal total cost for each input’s optimal solution; for example, the red dashed line marks the optimal cost of the blue line’s input $\mathbf{x}_{1}$ and its optimal solution $\mathbf{y}_{1}^{*}$.
@@ IV-C2 MSR Problem
Similarly, in Fig. 9, the vertical axis shows the total transmission rate, which we aim to maximize. Other settings of Fig. 9 follow the same format as Fig. 8. Since the high-quality solution space of individual inputs to this problem occupies a large area of the feasible solution space, the initial noise solution has performance close to the optimal solution, such as $\mathbf{x}_{2}$ in Fig. 9. In addition, we observe that MSR converges faster than CO, which aligns with the fact that MSR, as a convex optimization problem, has a lower complexity than the MINLP required for CO.
@@ V Conclusion
In this work, we bridge the theoretical gaps in GDM as an emerging network optimizer and provide further empirical guidance. This novel network optimization approach holds the potential to deliver significant improvements in both efficiency and quality for IoT networks. Our study has provided a concise theoretical demonstration of the inherent advantages of generative models over discriminative models in optimization solving. Specifically, GDMs offer a promising approach by learning the underlying high-quality solution distribution of feasible space and enabling repeated sampling from this distribution, which improves their ability to handle complex solution spaces and achieve optimal results. By capturing global solution structures rather than directly mapping inputs to outputs, GDMs address the inherent limitations of discriminative models. Our work has leveraged these strengths of GDMs, demonstrating their capacity to not only enhance the flexibility and accuracy of network optimization but also streamline the optimization process across alternative scenarios.